# Semiparametric robust estimation of population location

### v1

@@ Abstract
Real‑world measurements often comprise a dominant signal contaminated by a noisy background. Robustly estimating the dominant signal in practice has been a fundamental statistical problem. Classically, mixture models have been used to cluster the heterogeneous population into homogeneous components. Modeling such data with fully parametric models risks bias under misspecification, while fully nonparametric approaches can dissipate power and computational resources. We propose a middle path: a semiparametric method that models only the dominant component parametrically and leaves the background completely nonparametric, yet remains computationally scalable and statistically robust. So instead of outlier downweighting, traditionally done in robust statistics literature, we maximize the observed likelihood such that the noisy background is absorbed by the nonparametric component. Computationally, we propose a new approximate FFT-accelerated likelihood maximization algorithm. Empirically, this FFT plug‑in achieves order‑of‑magnitude speedups over vanilla weighted EM while preserving statistical accuracy and large sample properties.
@@ 1 Introduction
Finite mixture models have been a standard tool for modeling heterogeneous populations, discovering subpopulations, and performing model‑based clustering. Classical formulations impose a common parametric family on all components (e.g., Gaussian mixtures) and rely on identifiability assumptions or parsimonious parameterizations to render the estimation problem well‑posed [9, 4, 1, 8]. The EM algorithm is the workhorse in practice [32, 24, 10, 13]. Yet the assumptions that make parametric mixtures tractable—common family, structured covariances, or a fixed number of components—are almost never aligned with actual noisy datasets. In many applications the investigator primarily cares about the most dominant signal peak (its location and mass), while the remainder of the distribution is a complicated amalgam of weaker modes, tails, or background phenomena that defy clear parametric specification. For such tasks, fitting a fully parametric KK‑component model can be brittle: identifiability hinges on modeling all components correctly, and estimation can deteriorate due to over‑fitting or weak identifiability [31, 7].
@@
A parallel line of work relaxes parametric assumptions on one component. In two‑component mixtures where one distribution is known and the other is nonparametric (sometimes with symmetry or shape constraints), there are strong identifiability and estimation results for the mixing proportion and for features of the unknown component; see, e.g., [4] and [25]. In machine learning, the related mixture‑proportion estimation problem has been studied in semi‑supervised novelty detection and positive‑unlabeled learning, with estimators and rates under separability or irreducibility conditions [3, 27, 26]. These strands underscore a theme: if the inferential target is a weight or a single component, strict distributional modeling of the remainder can be avoided.
@@
Our paper embraces this philosophy and pushes it in a direction that is distinct from classical robust statistics. Robust methods à la Huber and Hampel downweight or reweight observations to resist the effect of outliers—with M‑estimators, influence‑function design, or divergence‑based criteria—but still typically posit a parametric data‑generating model and optimize a single loss over all data. Robust clustering extensions either replace Gaussian errors by heavier‑tailed $t$ components or trim a fraction of points to curb spurious solutions. While powerful for outlier resistance, these approaches differ in motivation and technique from our objective of extracting the dominant peak and its mass without modeling the rest [18, 16, 2].
@@
We consider a semiparametric “contamination” mixture $f(y)\;=\;\pi_{0}\,f_{0}(y;\theta)\;+\;(1-\pi_{0})\,f_{\mathrm{bg}}(y),\qquad\pi_{0}\in(0,1),$ in which $f_{0}(\cdot;\theta)$ is a parametric family for a dominant component (e.g., a location or location–scale family) and the background density $f_{\mathrm{bg}}^{*}$ is left completely unrestricted. The inferential target is the dominant component’s location parameter $\theta$. We do not attempt full identifiability of a multi‑component parametric mixture; rather, our goal is targeted recovery of the dominant component under certain regularity. This formulation sits at the intersection of semiparametric mixtures “with one unknown component” and mixture‑proportion estimation [4, 3, 27, 25, 26].
@@
Algorithmically, we maximize a plug–in log‑likelihood where the unknown background $f_{\mathrm{bg}}^{*}$ is replaced by a kernel density estimator (KDE) computed once per EM iteration via an FFT‑accelerated convolution on a uniform grid with linear binning and zero padding [28, 29, 35, 36]. This reduces the nonparametric update from $O(nG)$ to $O(G\log G)$ for $G$ grid points, while off‑grid evaluations use linear interpolation whose error we bound by the usual $C^{2}$ truncation rate $O(\Delta^{2}/h^{3})$ (here $\Delta$ is the grid spacing and $h$ the bandwidth), so the FFT/binned criterion differs from the exact plug–in criterion by a uniform $o_{p}(1)$ term.
@@
On the statistical side, we establish (i) consistency of the FFT plug–in estimator $\hat{\theta}_{n}$ under (a) standard smoothness/undersmoothing for KDE, (b) a windowed body–tail control for the log‑likelihood, and (c) a local curvature (strong concavity) condition for the population objective around $\theta^{*}$; and (ii) a CLT for smooth functionals after strengthening the discretization to $\sqrt{n}\,\varepsilon_{n}\!\to\!0$, where $\varepsilon_{n}\!=\!\tfrac{\|K^{\prime\prime}\|_{\infty}}{4}\Delta_{n}^{2}/h_{n}^{3}$ is the binned‑FFT interpolation error. The proofs combine a uniform oracle LLN for the parametric part [34] with (a) uniform control of the FFT/KDE approximation on a growing body window and (b) crude but sufficient tail envelopes that shrink against $\log(1/h_{n})$. The resulting consistency and CLT coincide with “what one would expect” if the background density were known, modulo the small plug–in remainder.
@@
Classical robust statistics mitigates sensitivity by modifying the loss, bounding the influence function, or minimizing a robust divergence [19, 15, 23]. Our approach differs in these aspects: (1) we do not posit a global parametric model for all observations—only the dominant component is parametric, the rest is nonparametric; (2) the weights arise from posterior membership in a mixture rather than from a downweighting scheme engineered via an influence curve. In this sense our method is closer to the semiparametric mixture literature, except here the unknown component is neither known nor shape‑constrained, and we leverage a fast nonparametric plug‑in evaluated inside the EM algorithm.
@@ 2 Estimating a single dominant component
Let $\Theta^{*}$ denote the truth of $\Theta$, and $f_{1}^{*}$ denote the truth of $f_{1}$ then from [37] we have the following theorem
@@
Under standard regularity and boundedness conditions  [37] guarantee the identifiability of $\hat{\Theta}$ and one can only identify the component that is large enough. The component with very small weight cannot be identified. A concavity assumption further guarantees the uniqueness of the parameter estimate.
The algorithm for the likelihood maximization proposed in [37] requires evaluating $\tilde{f}^{\mathrm{exact}}_{1,n}(y;\omega)=\frac{1}{h_{n}}\sum_{j=1}^{n}\omega_{j}\,\mathbb{K}\!\left(\frac{y_{j}-y}{h_{n}}\right)$ at all $Y_{i}$. This leads to a computational complexity of $\mathcal{O}(n^{2})$ per iteration. In finite samples their algorithm can also suffer from role flips where the parametric block absorbs the tight minority while the KDE models the diffuse majority. To counter such issues, we propose the following ideas.
@@ 2.1 Fast weighted KDE via binning and FFT
At grid points $x_{m},$ we have $x_{m}=x_{0}+m\Delta$ for $m=0,\ldots,M-1$. Define the sampled kernel $k_{h}[r]:=K_{h_{n}}(r\Delta)$ for integer lags $r\in\mathbb{Z}$. The binned KDE on the grid is the linear convolution
@@
In this FFT implementation due to first binning the data onto a grid, convolving these binned counts with the rescaled kernel $K_{h_{n}}$ on the grid via FFT, and recovering $\tilde{f}^{\mathrm{FFT}}_{1,n}(y)$ for arbitrary $y$ by linear interpolation, we encounter approximation error in two stages. A second–order Taylor expansion shows that both the binning step and the interpolation step introduce at most a $O(\Delta_{n}^{2}/h_{n}^{3})$ error per kernel term. Summing over the weights and using $\sum_{i}\hat{\omega}_{i}=1$ leads to a uniform bound on the discrepancy between the exact weighted KDE $\tilde{f}^{\mathrm{exact}}_{1,n}$ and its FFT approximation $\tilde{f}^{\mathrm{FFT}}_{1,n}$ over both the sample points $Y_{i}$ and the whole grid interval.
@@
$\max_{1\leq i\leq n}\big|\tilde{f}^{\mathrm{FFT}}_{1,n}(Y_{i})-\tilde{f}^{\mathrm{exact}}_{1,n}(Y_{i})\big|\ \leq\ \frac{1}{4}\,\|K^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
@@
$\sup_{y\in[x_{0},x_{M-1}]}\big|\tilde{f}^{\mathrm{FFT}}_{1,n}(y)-\tilde{f}^{\mathrm{exact}}_{1,n}(y)\big|\ \leq\ \frac{1}{4}\,\|K^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
@@ 2.3 Estimation algorithm
$\tilde{f}^{\mathrm{FFT},(t+1)}_{1,n}(x_{m})=\sum_{\ell=0}^{M-1}c^{(t)}[\ell]\;K_{h_{n}}(x_{m}-x_{\ell}),\qquad m=0,\ldots,M-1.$
@@
Data $\{Y_{i}\}_{i=1}^{n}$; kernel $\mathbb{K}$; bandwidth $h$; FFT grid $\{x_{j}\}_{j=0}^{M-1}$; tolerance tol; maximum iterations maxit. Randomly split the data by assigning $r_{i}^{(0)}\sim Ber(0.5).$ Initialize $\pi_{0}^{(0)},\,\theta^{(0)}$ from the M-step in Proposition 1
@@
  Background update: Normalise weights $\omega_{i}^{(t)}\leftarrow s_{i}^{(t)}\big/\sum_{k=1}^{n}s_{k}^{(t)}$; bin onto $\{x_{j}\}$ and convolve with $K_{h}$ via FFT to get $\{\tilde{f}_{1}^{(t)}(x_{j})\}_{j=0}^{M-1}$
@@ 3 Estimating more than 2 components
Let $(\hat{\theta}_{j},\hat{\alpha}_{j})$, $j<\ell$, denote the stage–(1,,1)($\dots,\ell-1$) estimators, and let $\hat{g}$ be an estimator of $g$ (for instance, the fitted mixture density from earlier stages). Define the unnormalised residual weights
$\widehat{v}_{i}^{(\ell)}:=1-\sum_{j<\ell}\hat{\alpha}_{j}\,\frac{f_{0}(Y_{i};\hat{\theta}_{j})}{\hat{g}(Y_{i})},\qquad i=1,\dots,n,$
and assume $\widehat{v}_{i}^{(\ell)}\geq 0$ a.s. for all $i$ and large $n$. Normalise to obtain probability weights
@@
$\sqrt{n}\big(\hat{\eta}_{\ell,n}-\eta_{\ell}^{*}\big)\ \rightsquigarrow\ \mathcal{N}\big(0,J^{(\ell)}(\eta_{\ell}^{*})^{-1}\big).$
In particular, the marginal distribution of $\sqrt{n}(\hat{\theta}_{\ell,n}-\theta_{\ell}^{*})$ is normal with covariance given by the $(\theta,\theta)$–block of $J^{(\ell)}(\eta_{\ell}^{*})^{-1}$.
@@
1:Data $\{Y_{i}\}_{i=1}^{n}$; stage $\ell\geq 1$; previous components $(\hat{\alpha}_{j},\hat{\theta}_{j})_{j<\ell}$ (empty if $\ell=1$); kernel $K$ with bandwidth $h_{n}$; FFT grid $\{x_{j}\}_{j=0}^{M-1}$ (spacing $\Delta$); tolerance tol; maximum iterations maxit.
@@
8:Initialisation
9:Choose $\theta_{\ell}^{(0)}$ (e.g. robust centre of $\{Y_{i}\}$ w.r.t. $\hat{w}_{i}^{(\ell)}$) and $\alpha_{\ell}^{(\ell,0)}\in(0,1)$.
10:Set $r_{i}^{(\ell,0)}\leftarrow s_{i}^{(\ell,0)}\leftarrow\tfrac{1}{2}$, $t\leftarrow 0$, $\tilde{\mathcal{L}}_{\ell}^{(0)}\leftarrow$-$\infty$.
@@
14:  Bin $\{(Y_{i},\omega_{i}^{(\ell,t)})\}$ on $\{x_{j}\}$ and convolve with $K_{h_{n}}$ via FFT to get $\{\tilde{f}_{\ell}^{(\mathrm{FFT},t)}(x_{j})\}$.

### v2

@@ Abstract
Real‑world measurements often comprise a dominant signal contaminated by a noisy background. Robustly estimating the dominant signal in practice has been a fundamental statistical problem. Classically, mixture models have been used to cluster the heterogeneous population into homogeneous components. Modeling such data with fully parametric models risks bias under misspecification, while fully nonparametric approaches can dissipate power and computational resources. We propose a middle path: a semiparametric method that models only the dominant component parametrically and leaves the background completely nonparametric, yet remains computationally scalable and statistically robust. So instead of outlier downweighting, traditionally done in robust statistics literature, we maximize the observed likelihood such that the noisy background is absorbed by the nonparametric component. Computationally, we propose a new FFT-accelerated approximate likelihood maximization algorithm. Empirically, this FFT plug‑in achieves order‑of‑magnitude speedups over vanilla weighted EM while preserving statistical accuracy and large sample properties.
@@ 1 Introduction
Finite mixture models have been a standard tool for modeling heterogeneous populations, discovering subpopulations, and performing model‑based clustering. Classical formulations impose a common parametric family on all components (e.g., Gaussian mixtures) and rely on identifiability assumptions or parsimonious parameterizations to render the estimation problem well‑posed [Bordes2006, ArcidiaconoJones2003, Dasgupta1999, DebSahaGuntuboyinaSen2022]. The EM algorithm is the workhorse in practice [Titterington1985, McLachlanPeel2000, Dempster1977, FraleyRaftery2002]. Yet the assumptions that make parametric mixtures tractable—common family, structured covariances, or a fixed number of components—are almost never aligned with actual noisy datasets. In many applications the investigator primarily cares about the most dominant signal peak (its location and mass), while the remainder of the distribution is a complicated amalgam of weaker modes, tails, or background phenomena that defy clear parametric specification. For such tasks, fitting a fully parametric KK‑component model can be brittle: identifiability hinges on modeling all components correctly, and estimation can deteriorate due to over‑fitting or weak identifiability [Teicher1963, Chen1995].
@@
A parallel line of work relaxes parametric assumptions on one component. In two‑component mixtures where one distribution is known and the other is nonparametric (sometimes with symmetry or shape constraints), there are strong identifiability and estimation results for the mixing proportion and for features of the unknown component; see, e.g., [Bordes2006] and [patra2016estimation]. In machine learning, the related mixture‑proportion estimation problem has been studied in semi‑supervised novelty detection and positive‑unlabeled learning, with estimators and rates under separability or irreducibility conditions [BlanchardLeeScott2010, Scott2015, RamaswamyScottTewari2016]. These strands underscore a theme: if the inferential target is a weight or a single component, strict distributional modeling of the remainder can be avoided.
@@
Our paper embraces this philosophy and pushes it in a direction that is distinct from classical robust statistics. Robust methods à la Huber and Hampel downweight or reweight observations to resist the effect of outliers—with M‑estimators, influence‑function design, or divergence‑based criteria—but still typically posit a parametric data‑generating model and optimize a single loss over all data. Robust clustering extensions either replace Gaussian errors by heavier‑tailed $t$ components or trim a fraction of points to curb spurious solutions. While powerful for outlier resistance, these approaches differ in motivation and technique from our objective of extracting the dominant peak and its mass without modeling the rest [Huber1964, Hampel1986, Basu1998].
@@
We consider a semiparametric “contamination” mixture $f(y)\;=\;\pi_{0}\,f_{0}(y;\theta)+(1-\pi_{0})\,f_{\mathrm{bg}}(y),\qquad\pi_{0}\in(0,1),$ in which $f_{0}(\cdot;\theta)$ is a parametric family for a dominant component (e.g., a location or location–scale family) and the background density $f_{\mathrm{bg}}$ is left completely unrestricted. The inferential target is the dominant component’s location parameter $\theta$. We do not attempt full identifiability of a multi‑component parametric mixture; rather, our goal is targeted recovery of the dominant component under certain regularity. This formulation sits at the intersection of semiparametric mixtures “with one unknown component” and mixture‑proportion estimation [Bordes2006, BlanchardLeeScott2010, Scott2015, patra2016estimation, RamaswamyScottTewari2016].
@@
Algorithmically, we maximize a plug–in log‑likelihood where the unknown background $f_{\mathrm{bg}}$ is replaced by a kernel density estimator (KDE) computed once per EM iteration via an FFT‑accelerated convolution on a uniform grid with linear binning and zero padding [Silverman1982, Silverman1986, Wand1994, WandJones1995]. This reduces the nonparametric update from $O(nG)$ to $O(G\log G)$ for $G$ grid points, while off‑grid evaluations use linear interpolation whose error we bound by the usual $C^{2}$ truncation rate $O(\Delta^{2}/h^{3})$ (here $\Delta$ is the grid spacing and $h$ the bandwidth), so the FFT/binned criterion differs from the exact plug–in criterion by a uniform $o_{p}(1)$ term.
@@
On the statistical side, we establish (i) consistency of the FFT plug–in estimator $\hat{\theta}_{n}$ under (a) standard smoothness/undersmoothing for KDE, (b) a windowed body–tail control for the log‑likelihood, and (c) a local curvature (strong concavity) condition for the population objective around $\theta^{*}$; and (ii) a CLT for smooth functionals after strengthening the discretization to $\sqrt{n}\,\varepsilon_{n}\!\to\!0$, where $\varepsilon_{n}\!=\!\tfrac{\|\mathbb{K}^{\prime\prime}\|_{\infty}}{4}\Delta_{n}^{2}/h_{n}^{3}$ is the binned‑FFT interpolation error. The proofs combine a uniform oracle LLN for the parametric part [vaart1998] with (a) uniform control of the FFT/KDE approximation on a growing body window and (b) crude but sufficient tail envelopes that shrink against $\log(1/h_{n})$. The resulting consistency and CLT coincide with “what one would expect” if the background density were known, modulo the small plug–in remainder.
@@
Classical robust statistics mitigates sensitivity by modifying the loss, bounding the influence function, or minimizing a robust divergence [Huber1981, Hampel1974, Maronna2019]. Our approach differs in these aspects: (1) we do not posit a global parametric model for all observations—only the dominant component is parametric, the rest is nonparametric; (2) the weights arise from posterior membership in a mixture rather than from a downweighting scheme engineered via an influence curve. In this sense our method is closer to the semiparametric mixture literature, except here the unknown component is neither known nor shape‑constrained, and we leverage a fast nonparametric plug‑in evaluated inside the EM algorithm.
@@ 2 Estimating a single dominant component
Let $\Theta^{*}$ denote the truth of $\Theta$, and $f_{1}^{*}$ denote the truth of $f_{1}$, then from [You2024sequential] we have the following theorem
@@
Under standard regularity and boundedness conditions  [You2024sequential] guarantee the identifiability of $\hat{\Theta}$ and one can only identify the component that is large enough. The component with very small weight cannot be identified. A concavity assumption further guarantees the uniqueness of the parameter estimate.
The algorithm for the likelihood maximization proposed in [You2024sequential] requires evaluating $\tilde{f}^{\mathrm{exact}}_{1,n}(y;\omega)=\frac{1}{h_{n}}\sum_{j=1}^{n}\omega_{j}\,\mathbb{K}\!\left(\frac{y_{j}-y}{h_{n}}\right)$ at all $Y_{i}$. This leads to a computational complexity of $\mathcal{O}(n^{2})$ per iteration. In finite samples their algorithm can also suffer from role flips where the parametric block absorbs the tight minority while the KDE models the diffuse majority. To counter such issues, we propose the following ideas.
@@ 2.1 Fast weighted KDE via binning and FFT
At grid points $x_{m},$ we have $x_{m}=x_{0}+m\Delta$ for $m=0,\ldots,M-1$. Define the sampled kernel $k_{h}[r]:=\mathbb{K}_{h_{n}}(r\Delta)$ for integer lags $r\in\mathbb{Z}$. The binned KDE on the grid is the linear convolution
@@
In this FFT implementation due to first binning the data onto a grid, convolving these binned counts with the rescaled kernel $\mathbb{K}_{h_{n}}$ on the grid via FFT, and recovering $\tilde{f}^{\mathrm{FFT}}_{1,n}(y)$ for arbitrary $y$ by linear interpolation, we encounter approximation error in two stages. A second–order Taylor expansion shows that both the binning step and the interpolation step introduce at most a $O(\Delta_{n}^{2}/h_{n}^{3})$ error per kernel term. Summing over the weights and using $\sum_{i}\hat{\omega}_{i}=1$ leads to a uniform bound on the discrepancy between the exact weighted KDE $\tilde{f}^{\mathrm{exact}}_{1,n}$ and its FFT approximation $\tilde{f}^{\mathrm{FFT}}_{1,n}$ over both the sample points $Y_{i}$ and the whole grid interval.
@@
$\max_{1\leq i\leq n}\big|\tilde{f}^{\mathrm{FFT}}_{1,n}(Y_{i})-\tilde{f}^{\mathrm{exact}}_{1,n}(Y_{i})\big|\ \leq\ \frac{1}{4}\,\|\mathbb{K}^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
@@
$\sup_{y\in[x_{0},x_{M-1}]}\big|\tilde{f}^{\mathrm{FFT}}_{1,n}(y)-\tilde{f}^{\mathrm{exact}}_{1,n}(y)\big|\ \leq\ \frac{1}{4}\,\|\mathbb{K}^{\prime\prime}\|_{\infty}\,\frac{\Delta_{n}^{2}}{h_{n}^{3}}.$
@@ 2.3 Estimation algorithm
$\tilde{f}^{\mathrm{FFT},(t+1)}_{1,n}(x_{m})=\sum_{\ell=0}^{M-1}c^{(t)}[\ell]\;\mathbb{K}_{h_{n}}(x_{m}-x_{\ell}),\qquad m=0,\ldots,M-1.$
@@
Data $\{Y_{i}\}_{i=1}^{n}$; kernel $\mathbb{K}$; bandwidth $h_{n}$; FFT grid $\{x_{j}\}_{j=0}^{M-1}$; tolerance tol; maximum iterations maxit. Randomly split the data by assigning $r_{i}^{(0)}\sim Ber(0.5).$ Initialize $\pi_{0}^{(0)},\,\theta^{(0)}$ from the M-step in Proposition 1
@@
  Background update: Normalise weights $\omega_{i}^{(t)}\leftarrow s_{i}^{(t)}\big/\sum_{k=1}^{n}s_{k}^{(t)}$; bin onto $\{x_{j}\}$ and convolve with $\mathbb{K}_{h_{n}}$ via FFT to get $\{\tilde{f}_{1}^{(t)}(x_{j})\}_{j=0}^{M-1}$
@@ 3 Estimating more than 2 components
Let $(\hat{\theta}_{j},\hat{\alpha}_{j})$, $j<\ell$, denote the stage–(1,,1)($\dots,\ell-1$) estimators, and let $\hat{f}$ be an estimator of $f$ (for instance, the fitted mixture density from earlier stages). Define the unnormalised residual weights
$\widehat{v}_{i}^{(\ell)}:=1-\sum_{j<\ell}\hat{\alpha}_{j}\,\frac{f_{0}(Y_{i};\hat{\theta}_{j})}{\hat{f}(Y_{i})},\qquad i=1,\dots,n,$
and assume $\widehat{v}_{i}^{(\ell)}\geq 0$ a.s. for all $i$ and large $n$. Normalize to obtain probability weights
@@
As before, denote $\mathbf{\theta}^{*}=(\theta_{1}^{*},\cdots,\theta_{J}^{*})$ as the truth of $\mathbf{\theta}$. Then the parameter estimates $\hat{\theta}_{1},\cdots,\hat{\theta}_{J}$ enjoy the following statistical properties.
@@
$\sqrt{n}\big(\hat{\eta}_{\ell,n}-\eta_{\ell}^{*}\big)\ \rightsquigarrow\ \mathcal{N}\big(0,I^{(\ell)}(\eta_{\ell}^{*})^{-1}\big).$
In particular, the marginal distribution of $\sqrt{n}(\hat{\theta}_{\ell,n}-\theta_{\ell}^{*})$ is normal with covariance given by the $(\theta,\theta)$–block of $I^{(\ell)}(\eta_{\ell}^{*})^{-1}$.
@@
1:Data $\{Y_{i}\}_{i=1}^{n}$; stage $\ell\geq 1$; previous components $(\hat{\alpha}_{j},\hat{\theta}_{j})_{j<\ell}$ (empty if $\ell=1$); kernel $\mathbb{K}$ with bandwidth $h_{n}$; FFT grid $\{x_{j}\}_{j=0}^{M-1}$ (spacing $\Delta$); tolerance tol; maximum iterations maxit.
@@
8:Initialization
9:Choose $\theta_{\ell}^{(0)}$ and $\alpha_{\ell}^{(\ell,0)}\in(0,1)$.
10:Set $r_{i}^{(\ell,0)}\leftarrow s_{i}^{(\ell,0)}\leftarrow\tfrac{1}{2}$, $t\leftarrow 0$.
@@
14:  Bin $\{(Y_{i},\omega_{i}^{(\ell,t)})\}$ on $\{x_{j}\}$ and convolve with $\mathbb{K}_{h_{n}}$ via FFT to get $\{\tilde{f}_{\ell}^{(\mathrm{FFT},t)}(x_{j})\}$.